5. Consider a process restringing tennis rackets. The process starts with a unit cost of $10 for the first unit—that is, c(1) = 10—and a learning rate of LR = 0.9. What will be the unit cost for the 35th unit?
Answer #1

Answer = 5.83

Explanation:

We will use the logarithmic approach:

C​​​​​​N = C​​​​​​​​​​​​1 N​​​​​​b

C​​​​​​N = Cost for Nth unit

C1 = Cost for the first unit

b = (log of the learning rate)/(log 2) = slope of the learning curve

Required = cost if 35th unit

= 10 × 35log0.9/log2

= 10 × 0.583

= 5.83